# Varying primordial state fractions in exo- and endothermic SIDM simulations of Milky Way-mass haloes

### v1

@@ Abstract
Self-interacting dark matter (SIDM) is increasingly studied as a potential solution to small-scale discrepancies between simulations of cold dark matter (CDM) and observations. We examine a physically motivated two-state SIDM model with both elastic and inelastic scatterings. In particular, endothermic, exothermic, and elastic scattering occur with equal probability at high relative velocities ($v_{\rm rel}\gtrsim 400\leavevmode\nobreak\ {\rm km/s})$. In a suite of cosmological zoom-in simulation of Milky Way-size haloes, we vary the primordial state fractions to understand the impact of inelastic dark matter self-interactions on halo structure and evolution. In particular, we test how the initial conditions impact the present-day properties of dark matter haloes. Depending on the primordial state fraction, scattering reactions will be dominated by either exothermic or endothermic effects for high and low initial excited state fractions respectively. We find that increasing the initial excited fraction reduces the mass of the main halo, as well as the number of subhaloes on all mass scales. The main haloes are cored, with lower inner densities and higher outer densities compared with CDM. Additionally, we find that the shape of the main halo becomes more spherical the higher the initial excited state fraction is. Finally, we show that the number of satellites steadily decreases with initial excited state fraction across all satellite masses.
@@ 1 Introduction
[…]
Although baryonic models may mitigate some of the small-scale discrepancies, it is difficult to resolve them all simultaneously. Since DM is also not well understood, there is significant interest in modifications to CDM that alleviate its small-scale problems while maintaining its large-scale accuracy (Cyr-Racine et al., 2016; Vogelsberger et al., 2016). In addition, it is also difficult to predict how effects from baryonic physics and alternative DM combine. Alternative DM models can alter star formation and gas distribution (Shen et al., 2024) and the stellar density profile (Vogelsberger et al., 2014).
@@
The parameter space of SIDM still remains largely unconstrained. Viable constant cross sections per mass ($\sigma_{\rm T}/m$) do not significantly alter the small-scale structure of galaxies (Zavala et al., 2013; Zeng et al., 2022; Silverman et al., 2023), so recent models have introduced velocity dependence such that the cross section depends on the relative velocity of the scattering particles (Vogelsberger et al., 2012; Meshveliani et al., 2022). SIDM particles also occur in many DM models, including atomic DM (Cline et al., 2014; Cyr-Racine & Sigurdson, 2013; Boddy et al., 2016), composite DM (Frandsen et al., 2011; Boddy et al., 2014; Chu et al., 2020), strongly interacting massive particles (Hochberg et al., 2014, 2015; Bernal et al., 2015; Choi et al., 2017; Hochberg et al., 2018), charged DM (McDermott et al., 2011; Liu et al., 2019; Dvorkin et al., 2019), mirror DM (Mohapatra et al., 2002; Foot, 2008; An et al., 2010; Chacko et al., 2021), and secluded DM (Boehm & Fayet, 2004; Pospelov et al., 2008; Feng & Kumar, 2008).
Additionally, SIDM scattering events can be elastic or inelastic. While many models initially explored elastic SIDM (Vogelsberger & Zavala, 2013; Rocha et al., 2013; Zavala et al., 2013; Elbert et al., 2015; Andrade et al., 2022), models including inelastic collisions are rising in interest due to their ability to alter the energy within the DM particles (Medvedev, 2010b, a, 2014a, 2014b; Schutz & Slatyer, 2015; Blennow et al., 2017; Zhang, 2017; Vogelsberger et al., 2019; Shen et al., 2021, 2022; O’Neil et al., 2023) and have important cosmological consequences (e.g. Fan et al., 2013; Foot, 2014; Vogelsberger et al., 2019; Xiao et al., 2021; Roy et al., 2023). One way to provide the mechanism for energy transfer during inelastic reactions is for the DM particles to exist in multiple states (e.g. Schutz & Slatyer, 2015; Vogelsberger et al., 2019; O’Neil et al., 2023). Other mechanisms, like decay (e.g. Wang et al., 2014), could have similar effects, but the multi-state model provides a large amount of flexibility and is what we focus on in this paper. The energy difference between the states can then be exchanged with the kinetic energy of the particles.
@@
In addition to the parameters of the SIDM model, it is also necessary to determine the initial abundance of states in the early universe, which depends on the cosmological history of the DM model. Previous work (e.g. Vogelsberger et al., 2019; O’Neil et al., 2023) has been restricted to initialising the DM particles entirely in one state or split evenly between the two states. The initial abundance of each state has marked effects on the resulting SIDM behaviour. If the particles are initially all in the ground state, for example, a model that is primarily exothermic will be indistinguishable from CDM. Generically, it is expected that the ground state is energetically favoured by the high-density environment of the early universe (Medvedev, 2001, 2010b, 2014b). Depending on the processes that determine the initial abundances, for example freeze-in or resonant annihilation, it is possible to obtain varying amounts of DM in the ground and excited states (Brahma et al., 2023).
@@ 2 Methods
Figure 1: The cross-sections as a function of relative particle velocity for each scattering reaction in the endothermic SIDM model. The top axis shows the halo $M_{200,\rm{mean}}$ with circular velocity $v$ at $R_{200,\rm{mean}}$. Elastic scattering between same state particles is very unlikely, while the elastic scattering between different state particles is similar in likelihood to down- and up-scattering. Down-scattering is most likely at lower relative velocities, and there is a velocity threshold for up-scattering to ensure that particles have enough kinetic energy to transform to the energy difference between the ground and excited states.
@@ 2.2 SIDM model
In this paper, we explore a DM model where the DM particles are able to self-interact with small cross-sections in addition to gravitational interactions. The particles scatter only a few times during the formation of the halo, so the DM remains largely collisionless. We employ the endothermic model introduced in O’Neil et al. (2023) based on the theory described in Schutz & Slatyer (2015) and Vogelsberger et al. (2019). In this framework, DM particles have two states, and the energy difference between the two states is much less than the ground-state energy. The model is fully described with four parameters: the ground state mass $m_{\chi^{1}}$, the mass splitting between the ground and excited states $\delta$, the dark force mediator mass $m_{\phi}$, and the dark coupling constant $\alpha$. The particles can scatter elastically, where there is no state change, or inelastically, where the particles change states. There are five ways for particles to scatter in our model:
@@
The values for the model we adopt in this work are $m_{\chi^{1}}=2.3\>$MeV, $\delta=0.48\>$eV, $m_{\phi}=7.3\>$MeV, and $\alpha=0.17$. These values lead to a significant probability of both up- and down-scattering in addition to elastic scattering between ground and excited state particles. Elastic scattering between either two ground state or two excited state particles is suppressed though does happen in small amounts. Figure 1 shows the cross section per unit mass ($\sigma/m_{\chi}$) for each type of scattering. The cross sections are a function of relative velocity between the two particles. The conversion between halo mass and relative velocity is done by assuming an isotropic Maxwell-Boltzmann velocity distribution of DM particles with one-dimensional velocity dispersion $\sigma_{\rm 1d}\simeq 0.6\,V_{\rm vir}$ (Łokas & Mamon, 2001), where $V_{\rm vir}\equiv\sqrt{GM_{\rm 200,mean}/R_{\rm 200,mean}}$ is the virial velocity of the halo. The elastic scattering between ground state particles (purple line) coincides with the cross section for elastic scattering between excited state particles (green dashed line). The up-scattering reaction (orange line) is only possible when the relative velocity is high enough such that there is sufficient kinetic energy to convert to the energy difference between the ground and excited states. At lower velocities, the down-scattering reaction (red line) becomes more likely, which can give particles more kinetic energy.
@@
We follow the algorithm described in Schutz & Slatyer (2015) and used in Vogelsberger et al. (2019) to implement the SIDM model in the simulations, which we briefly summarise here. Whether a particle scatters or not, and which scattering occurs, depends on both the scattering cross sections and the local density of dark matter. The local density is determined by the 10 to 38 nearest neighbours smoothed by a cubic spline kernel function around its nearest neighbours. The cross-section at the relevant velocity between two particles is multiplied by the DM (simulation macroscopic) particle mass and the distance each particle would move in one time step to get the probability of scattering. The probabilities are appropriately adjusted by the kernel factor as determined by the distance between the two potential scattering particles. Whether a particle scatters with a nearby particle and which reaction occurs is determined by drawing a random number.
In the event that a scattering reaction occurs, an appropriate second particle is selected to scatter with. The timesteps are set to be smaller than the expected scattering frequency so that multiple scatters per particle per time step are rare. In the case that there are multiple scatters in a single time step, the second reaction is rejected and does not take place. Rejected scatters make up less than $1\%$ of the total expected scatters, and further decreasing the timestep to avoid these does not impact the results. The particles scatter isotropically, and any changes to the particles’ energies are done in the centre of momentum frame of the two particles.
@@
Figure 4: The projected DM density for particles in the ground (left) or excited (right) states in the simulations with 0, 20, 40, 60, 80 or 100% of the particles starting in the excited state) at $z=0$. Most partices are in the ground state, and the halo does not change much in appearance between the ground state projections and the total particle projections. The excited state particles occur mostly in the centre of the halo, where the velocities are high enough for up-scattering to occur. The excited state particles also follow the trend of decreased density as the initial excited state fraction increases.
@@ 3 Results
Our suite is made up of 11 simulations with the DM particles initially $0-100\%$, increasing in intervals of $10\%$, in the excited state. We refer to the initial excited state fraction as $\chi_{\rm init}^{2}$ and the initial ground state fraction as $\chi_{\rm init}^{1}$. In Figure 3, we show DM projections for several simulations spanning the range of $\chi_{\rm init}^{2}$. Each simulation has a similar virial radius, ranging from 370.58 kpc for $\chi^{2}_{\text{init}}=100\%$ to 401.27 kpc for CDM.
@@ 3.1 Effects on halo density profiles
The mass distribution within the haloes varies significantly. The SIDM model gives a cored inner density profile for all initial excited fractions, compared to the cuspy profile of CDM. Within the SIDM simulations, however, increasing $\chi^{2}_{\text{init}}$ leads to a lower core density and higher outer density. The half mass radius $r_{1/2}$ increases significantly with $\chi^{2}_{\text{init}}$, while the virial radius $r_{200,\text{mean}}$ decreases slightly. In Figure 5, we plot density profiles for the main halo at $z=0$ in the top panel. Each SIDM halo has a core due to the influence of elastic scattering. Simulations with higher initial excited state fraction $\chi^{2}_{\textrm{init}}$ have lower inner densities, varying overall by two orders of magnitude.
@@
Shen et al. (2021) simulated a DM model that dissipates a constant fraction of a particle’s kinetic energy during an interaction. They ran a suite of zoom-in simulations of dwarf to Milky Way-mass galaxies with different cross sections in addition to a fiducial CDM model, including the baryonic physics described by the FIRE-2 model (Hopkins et al., 2018). They found that the halo central density increases with cross section (and effectively the energy dissipation rate of DM). Similarly to the up-scattering in our work, the dissipation causes a particle to lose velocity and fall inwards, thus increasing the central density.
Exothermic reactions, on the other hand, decrease the central density, as explored in Vogelsberger et al. (2019). They included both elastic and exothermic reactions in $N$-body zoom-in simulations, but endothermic reactions were suppressed.
@@
An initially larger fraction of excited particles eventually provides heating of DM particles through downscattering, pushing more particles away from the halo center. This also leads to higher densities at $R_{200}$ for simulations with higher $\chi^{2}_{\textrm{init}}$, although the difference is less dramatic. The velocity boost also causes virial mass to decrease with $\chi^{2}_{\textrm{init}}$ as more particles are driven above escape velocity as shown in Figure 6.
@@
The velocity dispersions in the SIDM models reach an equilibrium at lower radii while the CDM peaks at about $10$ kpc. Because heat can transfer between DM particles in a self-interacting model, the velocity dispersion smooths out, The equilibrium velocity dispersion decreases with $\chi_{\rm init}^{2}$, and it is higher than the CDM case for low $\chi_{\rm init}^{2}$. This implies that there is heat transferred out of the halo when particles start in the excited state. Initial down-scattering pushes particles outwards and out of the halo, leaving the halo at a slightly lower mass and with less energy. In the low $\chi^{2}_{\rm init}$ simulations, on the other hand, up-scattering earlier in the halo’s evolution creates a deep gravitational well that binds particles within the halo. The mass of the $\chi_{\rm init}^{2}=0\%$ is $2.16\times 10^{12}\>{\,\rm M_{\odot}}$ and $\chi_{\rm init}^{2}=100\%$ is $1.70\>{\,\rm M_{\odot}}$, and the mass decreases monotonically between the two.
@@
[…]

### v2

@@ Abstract
Self-interacting dark matter (SIDM) is increasingly studied as a potential solution to small-scale discrepancies between simulations of cold dark matter (CDM) and observations. We examine a physically motivated two-state SIDM model with both elastic and inelastic scatterings. In particular, endothermic, exothermic, and elastic scattering have equal transfer cross sections at high relative velocities ($v_{\rm rel}\gtrsim 400\leavevmode\nobreak\ {\rm km/s})$. In a suite of cosmological zoom-in simulation of Milky Way-size haloes, we vary the primordial state fractions to understand the impact of inelastic dark matter self-interactions on halo structure and evolution. In particular, we test how the initial conditions impact the present-day properties of dark matter haloes. Depending on the primordial state fraction, scattering reactions will be dominated by either exothermic or endothermic effects for high and low initial excited state fractions respectively. We find that increasing the initial excited fraction reduces the mass of the main halo, as well as the number of subhaloes on all mass scales. The main haloes are cored, with lower inner densities and higher outer densities compared with CDM. Additionally, we find that the shape of the main halo becomes more spherical the higher the initial excited state fraction is. Finally, we show that the number of satellites steadily decreases with initial excited state fraction across all satellite masses.
@@ 1 Introduction
Many prescriptions for baryonic processes rely on subgrid models, where the relevant physics occurs at scales less than the resolution of the simulation. The effects of these subgrid models are sensitive to model parameters and can vary significantly and impact key comparisons between simulations and observations (Vogelsberger et al., 2020). The hydrodynamics solver for gas, e.g. using the moving-mesh scheme Arepo or the smoothed particle hydrodynamics scheme Gadget, can change the size of disks that are produced (Torrey et al., 2012). The star formation models are also not well understood. Altering the density threshold for star formation in gas influences whether cores are formed in dwarf galaxies Benítez-Llambay et al. (2019), altering the primary factor that determines the star formation rate, e.g. a density threshold or an H2-based model, changes the number of satellites by a factor of two (Munshi et al., 2019), and altering the star formation efficiency affects the morphology of galaxies (Agertz & Kravtsov, 2015). These parameters can thus be the difference between matching observations or producing unrealistic galaxies.
Although baryonic models may mitigate some of the small-scale discrepancies, it is difficult to resolve them all simultaneously. Since DM is also not well understood, there is significant interest in modifications to CDM that alleviate its small-scale problems while maintaining its large-scale accuracy (Cyr-Racine et al., 2016; Vogelsberger et al., 2016). In addition, it is also difficult to predict how effects from baryonic physics and alternative DM combine. Alternative DM models can alter star formation and gas distribution (Shen et al., 2024a) and the stellar density profile (Vogelsberger et al., 2014).
@@
The parameter space of SIDM still remains largely unconstrained. Viable constant transfer cross sections per mass ($\sigma_{\rm T}/m$) do not significantly alter the small-scale structure of galaxies (Zavala et al., 2013; Zeng et al., 2022; Silverman et al., 2023), so recent models have introduced velocity dependence such that the cross section depends on the relative velocity of the scattering particles (Vogelsberger et al., 2012; Meshveliani et al., 2023). SIDM particles also occur in many DM models, including atomic DM (Cline et al., 2014; Cyr-Racine & Sigurdson, 2013; Boddy et al., 2016), composite DM (Frandsen et al., 2011; Boddy et al., 2014; Chu et al., 2020), strongly interacting massive particles (Hochberg et al., 2014, 2015; Bernal et al., 2015; Choi et al., 2017; Hochberg et al., 2018), charged DM (McDermott et al., 2011; Liu et al., 2019; Dvorkin et al., 2019), mirror DM (Mohapatra et al., 2002; Foot, 2008; An et al., 2010; Chacko et al., 2021), and secluded DM (Boehm & Fayet, 2004; Pospelov et al., 2008; Feng & Kumar, 2008).
Additionally, SIDM scattering events can be elastic or inelastic. While many models initially explored elastic SIDM (Vogelsberger & Zavala, 2013; Rocha et al., 2013; Zavala et al., 2013; Elbert et al., 2015; Andrade et al., 2022), models including inelastic collisions are rising in interest due to their ability to alter the energy within the DM particles (Medvedev, 2010b, a, 2014a, 2014b; Schutz & Slatyer, 2015; Blennow et al., 2017; Zhang, 2017; Vogelsberger et al., 2019; Shen et al., 2021, 2024b; O’Neil et al., 2023) and have important cosmological consequences (e.g. Fan et al., 2013; Foot, 2014; Vogelsberger et al., 2019; Xiao et al., 2021; Roy et al., 2023). One way to provide the mechanism for energy transfer during inelastic reactions is for the DM particles to exist in multiple states (e.g. Schutz & Slatyer, 2015; Vogelsberger et al., 2019; O’Neil et al., 2023). Other mechanisms, like decay (e.g. Wang et al., 2014), could have similar effects, but the multi-state model provides a large amount of flexibility and is what we focus on in this paper. The energy difference between the states can then be exchanged with the kinetic energy of the particles.
@@
In addition to the parameters of the SIDM model, it is also necessary to determine the initial abundance of states in the early universe, which depends on the cosmological history of the DM model. Previous work (e.g. Vogelsberger et al., 2019; O’Neil et al., 2023) has been restricted to initialising the DM particles entirely in one state or split evenly between the two states. The initial abundance of each state has marked effects on the resulting SIDM behaviour. If the particles are initially all in the ground state, for example, a model that is primarily exothermic will be indistinguishable from CDM. Generically, it is expected that the ground state is energetically favoured by the high-density environment of the early universe (Medvedev, 2001, 2010b, 2014b). Depending on the processes that determine the initial abundances, for example freeze-in or resonant annihilation, it is possible to obtain varying amounts of DM in the ground and excited states (Brahma et al., 2024).
@@ 2 Methods
Figure 1: The transfer cross sections as a function of relative particle velocity for each scattering reaction in the endothermic SIDM model. The top axis shows the halo $M_{200,\rm{mean}}$ with circular velocity $v$ at $R_{200,\rm{mean}}$. Elastic scattering between same state particles is very unlikely, while the elastic scattering between different state particles is similar in likelihood to down- and up-scattering. Down-scattering is most likely at lower relative velocities, and there is a velocity threshold for up-scattering to ensure that particles have enough kinetic energy to transform to the energy difference between the ground and excited states.
@@ 2.2 SIDM model
In this paper, we explore a DM model where the DM particles are able to self-interact with small cross-sections in addition to gravitational interactions. The particles scatter only a few times during the formation of the halo, so the DM remains largely collisionless. We employ the endothermic model introduced in O’Neil et al. (2023) based on the theory described in Section 3.1 ofSchutz & Slatyer (2015). The code is the same as used in Vogelsberger et al. (2019), which modelled a similar two-state model with different cross sections. In this framework, DM particles have two states, and the energy difference between the two states is much less than the ground-state energy. The model is fully described with four parameters: the ground state mass $m_{\chi^{1}}$, the mass splitting between the ground and excited states $\delta$, the dark force mediator mass $m_{\phi}$, and the dark coupling constant $\alpha$. The particles can scatter elastically, where there is no state change, or inelastically, where the particles change states. There are five ways for particles to scatter in our model:
@@
The values for the model we adopt in this work are $m_{\chi^{1}}=2.3\>$MeV, $\delta=0.48\>$eV, $m_{\phi}=7.3\>$MeV, and $\alpha=0.17$. These values lead to a significant probability of both up- and down-scattering in addition to elastic scattering between ground and excited state particles. Elastic scattering between either two ground state or two excited state particles is suppressed though does happen in small amounts. While there have been some constraints placed on elastic scattering (see Adhikari et al., 2022, for review), inelastic models are less constrained. Additionally, these cross sections are high enough to be on the extreme end of what might be seen, both making the influences clear and testing the limits of the model.
Figure 1 shows the transfer cross section per unit mass ($\sigma_{T}/m_{\chi}$) for each type of scattering. The transfer cross sections are a function of relative velocity between the two particles. The conversion between halo mass and relative velocity is done by assuming an isotropic Maxwell-Boltzmann velocity distribution of DM particles with one-dimensional velocity dispersion $\sigma_{\rm 1d}\simeq 0.6\,V_{\rm vir}$ (Łokas & Mamon, 2001), where $V_{\rm vir}\equiv\sqrt{GM_{\rm 200,mean}/R_{\rm 200,mean}}$ is the virial velocity of the halo. The elastic scattering between ground state particles (purple line) coincides with the cross section for elastic scattering between excited state particles (green dashed line) at a value of $0.21\>{\rm cm}^{2}{\rm g}^{-1}$. The up-scattering reaction (orange line) is only possible when the relative velocity is high enough such that there is sufficient kinetic energy to convert to the energy difference between the ground and excited states. At lower velocities, the down-scattering reaction (red line) becomes more likely, scaling with a power law of -1, which can give particles more kinetic energy. Elastic scattering between a ground and an excited state particle is approximately constant at a value of $15\>{\rm cm}^{2}{\rm g}^{-1}$.
@@
We follow the algorithm described in Schutz & Slatyer (2015) and used in Vogelsberger et al. (2019) to implement the SIDM model in the simulations, which we briefly summarise here. Whether a particle scatters or not, and which scattering occurs, depends on both the scattering transfer cross sections and the local density of dark matter. The local density is determined by the 10 to 38 nearest neighbours smoothed by a cubic spline kernel function around its nearest neighbours. The cross-section at the relevant velocity between two particles is multiplied by the DM (simulation macroscopic) particle mass and the distance each particle would move in one time step to get the probability of scattering. The probabilities are appropriately adjusted by the kernel factor as determined by the distance between the two potential scattering particles. Whether a particle scatters with a nearby particle and which reaction occurs is determined by drawing a random number.
In the event that a scattering reaction occurs, an appropriate second particle is selected to scatter with. The timesteps are set to be smaller than the expected scattering frequency so that multiple scatters per particle per time step are rare. In the case that there are multiple scatters in a single time step, the second reaction is rejected and does not take place. This is because a particle’s position and velocity can be updated only once per time step. Rejected scatters make up less than $1\%$ of the total expected scatters, and further decreasing the time step to avoid these does not impact the results. We have tested this by running the same initial conditions with larger cross sections and varying the time step and rejection rates under 5 percent do not noticeable alter the particle distribution in the halo. The particles scatter isotropically, and any changes to the particles’ energies are done in the centre of momentum frame of the two particles.
@@
Figure 4: The projected DM density for particles in the ground (left) or excited (right) states in the simulations with 0, 20, 40, 60, 80 or 100% of the particles starting in the excited state at $z=0$. Most partices are in the ground state, and the halo does not change much in appearance between the ground state projections and the total particle projections. The excited state particles occur mostly in the centre of the halo, where the velocities are high enough for up-scattering to occur. The excited state particles also follow the trend of decreased density as the initial excited state fraction increases.
@@ 3 Results
Our suite is made up of 11 simulations with the DM particles initially $0-100\%$, increasing in intervals of $10\%$, in the excited state. We refer to the initial excited state fraction as $\chi_{\rm init}^{2}$ and the initial ground state fraction as $\chi_{\rm init}^{1}$. In Figure 2, we show the DM projection for our fiducial CDM simulation. In Figure 3, we show DM projections for all particles several simulations spanning the range of $\chi_{\rm init}^{2}$. We show the DM projections for excited and ground state particles separately in Figure 4. Each simulation has a similar virial radius, ranging from 370.58 kpc for $\chi^{2}_{\text{init}}=100\%$ to 401.27 kpc for CDM.
@@ 3.1 Effects on halo density profiles
The mass distribution within the haloes varies significantly. In Figure 5, we plot density profiles for the main halo at $z=0$ in the top panel. The SIDM model gives a cored inner density profile for all initial excited fractions, compared to the cuspy profile of CDM. This is similar to the effects of classical elastic SIDM models previously explored in literature (e.g. Zavala et al., 2013; Rocha et al., 2013; Elbert et al., 2015) and is due to thermalisation of DM cores from heat conduction.
@@
The two-state configuration of the SIDM model studied in this paper enables additional channels of heating or cooling in the dark sector. As shown in Figure 5, when $\chi^{2}_{\text{init}}$ is low, we find that the central density of the simulated Milky Way-mass halo is enhanced compared to CDM. This is similar to the gravothermal core-collapse studied in lower halo mass regimes and often in SIDM models with strong velocity-dependence of cross sections (e.g. Nadler et al., 2020; Sameie et al., 2020; Correa et al., 2022; Zeng et al., 2022). The up-scatterings of DM particles dissipate away their kinetic energy and can accelerate the gravothermal collapse of the halo (e.g. Essig et al., 2019). This is a common phenomenon found in previous explorations of dissipative SIDM (e.g. Essig et al., 2019; Shen et al., 2021; Roy et al., 2023). For example, Shen et al. (2021) simulated a DM model that dissipates a constant fraction of a particle’s kinetic energy during an interaction. They ran a suite of zoom-in simulations of dwarf to Milky Way-mass galaxies with different cross sections in addition to a fiducial CDM model, including the baryonic physics described by the FIRE-2 model (Hopkins et al., 2018). They found that the halo central density increases with cross section and effectively the energy dissipation rate of DM. Similar to the up-scattering in our work, the dissipation causes a particle to lose velocity and fall inwards, thus increasing the central density.
@@
However, increasing $\chi^{2}_{\text{init}}$ leads to a lower core density and higher outer density. The half mass radius $r_{1/2}$ increases significantly with $\chi^{2}_{\text{init}}$, while the virial radius $r_{200,\text{mean}}$ decreases slightly. As shown in the top panel of Figure 5, simulations with higher initial excited state fraction $\chi^{2}_{\textrm{init}}$ have lower inner densities, varying overall by two orders of magnitude. Exothermic reactions from down-scatterings dominate in these cases and decrease the central density, as previously explored in $N$-body zoom-in simulations (Vogelsberger et al., 2019). An initially larger fraction of excited particles eventually provides heating of DM particles through downscattering, pushing more particles away from the halo center. This also leads to higher densities at $R_{200}$ for simulations with higher $\chi^{2}_{\textrm{init}}$, although the difference is less dramatic. The velocity boost also causes virial mass ($M_{200}$) to decrease with $\chi^{2}_{\textrm{init}}$ as more particles are driven above escape velocity as shown in Figure 6.
@@
The gradients of the velocity dispersions in the SIDM models vanish at lower radii while the CDM peaks at about $10$ kpc. Because heat can transfer between DM particles in a self-interacting model, the velocity dispersion smooths out, The equilibrium velocity dispersion decreases with $\chi_{\rm init}^{2}$, and it is higher than the CDM case for low $\chi_{\rm init}^{2}$. This implies that there is heat transferred out of the halo when particles start in the excited state. Initial down-scattering pushes particles outwards and out of the halo, leaving the halo at a slightly lower mass and with less energy. In the low $\chi^{2}_{\rm init}$ simulations, on the other hand, up-scattering earlier in the halo’s evolution creates a deep gravitational well that binds particles within the halo. The mass of the $\chi_{\rm init}^{2}=0\%$ is $2.16\times 10^{12}\>{\,\rm M_{\odot}}$ and $\chi_{\rm init}^{2}=100\%$ is $1.70\>{\,\rm M_{\odot}}$, and the mass decreases monotonically between the two.
@@
In Figure 6 we show M200,mean and Vmax for the main halo in each simulation. $V_{\rm max}$ is the maximum circular velocity of the halo. As a result of dominant upscatterings, SIDM models with low $\chi_{\rm init}^{2}$ values produce more concentrated DM distributions and thus give larger $V_{\rm max}$ values. On the contrary, models with large $\chi_{\rm init}^{2}$ values feature substantial downscatterings of DM particles that create cored density profiles. The cores not only decrease the $V_{\rm max}$ values in these models but also lead to lower halo virial radius and masses (given the same spherical overdensity threshold).